OpenStudy (anonymous):

Write the equation of the line that passes through (–1, 4) with a slope of –1 in standard form.

OpenStudy (anonymous):

y=mx+c the slope is given so insert the "-1" into m.. you'll get y=-x+c to get the value of c(y-intercept) insert both value of x and y given in the question: 4=-(-1)+c 4= 1+c 3 = c c= 3 so the equation would be y=-x+3